You are long 20 gold futures contracts, established at an initial settle price of $1,512 per ounce, where each contract represents 100 troy ounces. Your initial margin to establish the position is $12,000 per contract and the maintenance margin is $11,200 per contract. Over the subsequent four trading days, gold settles at $1,506, $1,502, $1,507, and $1,517, respectively.

Compute the balance in your margin account at the end of each of the four trading days, and compute your total profit or loss at the end of the trading period. Assume that a margin call requires you to fund your account back to the initial margin requirement.
Answer #1

15 1 A F G H 2 6 4 Price Opening Balance Change in margin Balance Add Closing 5 1512 240000 240000 240000 1506 240000 -12000

D Margin Per contract =F5/20 =F6/20 =F7/20 =F8/20 =F9/20 4 Price Opening Balance Change in margin Balance Add Closing 5 1512